OpenStudy (anonymous):

Use the properties of logarithms to write the following expression as a sum or difference of simple logarithmic term. In(x4square root y)

OpenStudy (callisto):

Do you mean \[\ln x^4\sqrt{y}\] for your question?

OpenStudy (anonymous):

no 4v

OpenStudy (anonymous):

v4

OpenStudy (callisto):

\[\ln x \sqrt[4]{y}\] ?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

\[\ln x \sqrt[4]{y}\]\[=\ln x + \ln\sqrt[4]{y}\]\[=\ln x + \ln y^{1/4}\]\[=\ln x + (1/4)\ln y \]

OpenStudy (anonymous):

Thank you

OpenStudy (callisto):

Welcome, but do you understand how to do it (or why I did it like that) ?

OpenStudy (anonymous):

I always get confused of with those question the logs and lns? do u have any helpful points to help me with?

OpenStudy (callisto):

Some, perhaps \[\log x^n =nlogx\] \[\log xy =logx + logy\]

OpenStudy (callisto):

Those are for your question only and also you might find this helpful too \[log(a/b) = loga -logb\]

OpenStudy (anonymous):

are lns the same BUT the diffrent is ln=e and log=10 right??

OpenStudy (callisto):

Plus one more point \[\sqrt[n]{x} = x ^{(1/n)}\]So, \[\log \sqrt[n]{x} = \log x ^{(1/n)}= (1/n)\log x \]

OpenStudy (callisto):

Hmm.. these are the properties of logarithm, no matter what base it is , you can still apply it ln is natural log with base e, while log is the common log, with base 10 which means lne =1, while log10 =1

OpenStudy (callisto):

Hmm, does that make it better or worse?

OpenStudy (anonymous):

of course, thank you also, now i understand this part, but when should I use the e and ln + 10 and log when the question says what?

OpenStudy (callisto):

I think it depends on the situation. If you are like given e in the question, probably using ln is more convenient, or else just use the common log. Plus, sometimes the question itself may give you the log with other bases, like 2, and so on. So use the one you think it is the most convenient :)

OpenStudy (anonymous):

that what I hate about math... lol :) thank you. i got sick of the complex stuff ugh. thank u very much again

OpenStudy (callisto):

(You're welcome) x2 . Math could be fun with more practice :)

OpenStudy (anonymous):

that is as like as our recent philosophy of life.

OpenStudy (anonymous):

thank u

OpenStudy (callisto):

You're welcome again :)

OpenStudy (anonymous):

Also, how to answer this ln (8x-4)=ln2+lnx